You are going to value davids Doll Co. using the FCF model. After consulting various sources, you find that Lauryn's has a reported equity beta of 1.7, a debt-to-equity ratio of .6, and a tax rate of 21 percent. Assume a risk-free rate of 3 percent and a market risk premium of 7 percent. Davids Doll Co. had EBIT last year of $58 million, which is net of a depreciation expense of $5.8 million. In addition, David made $6.3 million in capital expenditures and increased net working capital by $2.5 million. Assume the FCF is expected to grow at a rate of 3 percent into perpetuity. What is the value of the firm?
Answer #1

Equity Beta = Beta * (1 + debt to equity ratio * (1 - debt to equity ratio))

1.7 = beta * (1 + .6*(1 - .6))

1.7 = beta * (1.24)

beta = 1.7/1.24

beta = 1.37

FCF of Davids Doll Co = EBIT * (1 - tax rate) + depreciation - capex - net working capital

= 58 * (1 - .21) + 5.8 - 6.3 - 2.5

= 42.82

required rate of return = risk free rate + beta * market premium

= 3% + 1.37 * 7%

= 12.6%

Value of firm = FCF * (1 + growth rate)/(required rate of return - growth rate)

= 42.82 * (1.03)/(.096)

= $459.42 Million